Elementary Functions and Calculus I Math 131 (Sec 42), Autumn 2004 Course Document

Lecturer: David Rule Office: Eckhart 16 Phone: (773) 702 6045 Email: rule@math.uchicago.edu Website: www.math.uchicago.edu/˜ rule/131/

Lectures: Monday, Wednesday & Friday, 11:30am-12:20pm, Room SS105.

Tutorials: Tuesday & Thursday, 12:00am-1:20pm Marina Danilevsky (mdanilev@uchicago.edu) Kent 103 Kevin Choi (kylchoi@uchicago.edu) Kent 106 Kwokfung Tang (kftang@uchiacgo.edu) Cobb 206

Office hours: Wednesday 12:30-1:30pm and Thursday 4-5pm

Text: Calculus, Varberg, Purcell & Rigdon, Chapters 1,2 & 3.

Syllabus: The 131-2-3 sequence aims to teach the basic concepts and techniques of rigorous mathematics and apply these to certain problems. Included in these aims is presenting you work in an understandable and coherent manner. The course 131 will cover the real number system, inequalities, absolute values, functions, limits and limit theorems, the derivative, the chain rule, higher-order derivatives and implicit differentiation.

Assessment: Your grade will be determined by two in-class mid-term tests, a final examination, and homework. I may occasionally give quizzes but these will not count towards your grade. Homework should be handed in on or before the given deadline. Late homework will be given zero credit unless an extension is given by me. Extensions and the like will only be given in exceptional circumstances (for example, sudden ill health) and should be requested before the deadline. The final constitutes 40% of your grade, the midterms 20% each, and homework totals 20%.

The Final Examination: The final examination is on Wednesday, 8th December from 10:30am to 12:30pm. It is the policy of the Department of Mathematics that the following rules apply to final exams in all undergraduate mathematics courses:

1. The final exam must occur at the time and place designated on the College Final Exam Schedule. In particular, no final examinations may be given during the tenth week of the quarter, except in the case of graduating seniors.

2. Any student who wishes to depart from the scheduled final exam time for the course must receive permission from Paul Sally (office is Ry 350, phone is 2-7388, email is sally@math.uchicago.edu). Instructors are not permitted to excuse students from the scheduled time of the final exam except in the cases of an Incomplete. The University examination schedule can be found at http://registrar.uchicago.edu/students/final-exams.html.

Miscellaneous Items: The use of calculators, both for homework and during examinations, is forbidden. The rule is a little redundant as you will find them next to useless.

David Rule